OpenStudy (anonymous):

Which ordered pair is not a solution of y = –12x – 3? (1 point) • (–3, 33) • (8, –98) • (–9, 105) • (3, –39)

OpenStudy (amorfide):

if you let a standard coordinate set be (x,y) then you know that from (-3,33) that x=-3, y=33 substitute into the linear equation given and see if it works out if not try the next set

OpenStudy (texaschic101):

your points are set up like this (x,y) so just sub your points into the equation and see which ones come out true. I will do one for you... y = -12x - 3 (-3,33)...x = -3 and y = 33 now sub 33 = -12(-3) - 3 33 = 36 - 3 33 = 33 (correct)...this one happens to be correct. However, you are looking for the one that is not correct...so just plug in your points and find the one that does not work

OpenStudy (texaschic101):

oops..typo...see which ones come out false because you are looking for the one that is not correct

OpenStudy (anonymous):

so its A